The average length of time for students to enroll in college is 6 hours and 10 minutes. Computerized enrolment is being tested. In a random sample of 180 students, it was found out that the average time for them to enroll is 2 hours and 45 minutes with a standard deviation of 20 minutes. At = 0.01, does this indicate that the average time for students to enroll using computerized enrolment is less than the regular enrolment?

1
Expert's answer
2022-05-23T12:18:46-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge370"

"H_a:\mu<370"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=179" degrees of freedom, and the critical value for a left-tailed test is "t_c =-2.34736."

The rejection region for this left-tailed test is "R = \{t:t< -2.34736\}."

The t-statistic is computed as follows:



"t=\dfrac{165-370}{20/\sqrt{180}}=-137.518"

Since it is observed that "t =-137.518<-2.34736=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for left-tailed "df=179" degrees of freedom, "t=-137.518" is "p=0," and since "p=0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than 190, at the "\alpha = 0.01" significance level.